In an ideal Otto cycle, air is compressed from 1.20 kg/m3 and 2.2 to 0.26 L, and the net work output of the cycle is 440 kJ/kg. The mean effective pressure (MEP) for this cycle is
1
Expert's answer
2021-07-19T15:12:22-0400

Mean effective pressure (MEP)

"MEP=\frac{w}{\alpha_2-\alpha_1}=\frac{wm}{V_1-V_2}"

MEP="\frac{\rho_1V_1w_1}{V_1-V_2}"

Put all value


"MEP=\frac{1.20\times2.2\times440}{2.2-0.26}=599\times10^3 pa"

"MEP=599\times10^{3}pa=599Kpa"